# Diffusion Models for SU(2) Lattice Gauge Theory in Two Dimensions

### v1

@@ I Introduction
The emergence of machine learning approaches offers promising avenues to address these challenges. Normalizing flows have been successfully applied to scalar field theories Albergo2019 and gauge theories Kanwar2020 ; Boyda2021 ; Abbott2023 , demonstrating the potential of generative models in lattice field theory. More recently, diffusion models—a class of generative models that learn to reverse a gradual noising process—have achieved remarkable success in image generation Ho2020 ; Song2021 and have been proposed for lattice gauge theory applications Wang:2023exq ; Wang:2023sry ; Zhu2024a ; Zhu2024b ; Aarts:2026zzr .
@@ IV.4 Training Procedure
The training procedure follows the standard DDPM approach, as illustrated in Fig. 3. For each batch of clean configurations $\varphi_{0}$:
@@
Figure 3: Flowchart of a single training step. The model learns to predict the noise $\epsilon_{\text{true}}$ that was added to clean configurations, enabling the reverse denoising process during sampling.
@@ IV.5 Sampling Algorithm
To generate new configurations, we reverse the diffusion process starting from pure noise, as illustrated in Fig. 4.
Figure 4: Sampling algorithm flowchart. Starting from Gaussian noise, the iterative denoising loop applies physics-conditioned noise prediction and gradually recovers a gauge field configuration. The final projection step ensures the output satisfies the SU(2) constraint.
@@ V.1 Validation at Training Coupling
Figure 5 shows the plaquette distribution for the training point. The distribution is well-centered on the exact value, with the sample mean nearly coinciding with it, confirming that the model has learned the correct equilibrium distribution at $(\beta_{0},8\times 8)$.
Figure 5: Plaquette distribution from 500 diffusion-generated configurations at $\beta=2.0$ on the $8\times 8$ training lattice. Red dashed line: exact value $I_{2}(\beta)/I_{1}(\beta)$; blue line: sample mean. The distribution is well-centered on the exact result.
@@ V.2 Physics-Conditioned Generation at Different Couplings
Figure 6 shows the average plaquette as a function of $\beta$. Near the training coupling, the model achieves excellent agreement: at $\beta=1.5$, $2.0$, and $2.5$ the biases are $|\Delta|\leq 0.0004$. At the boundaries of the explored coupling range, larger biases appear: $|\Delta|=0.017$ at $\beta=1.0$ and $|\Delta|=0.051$ at $\beta=3.0$, with intermediate values $|\Delta|\approx 0.01$$0.03$ at $\beta=3.5$ and $4.0$. This demonstrates that physics-conditioned rescaling captures the coupling dependence of the score function well near the training point, with accuracy degrading gradually at larger $|\beta-\beta_{0}|$.
Figure 6: Average plaquette $\langle P\rangle$ as a function of coupling $\beta$ for the $8\times 8$ training lattice. Blue circles: diffusion model results with $\sigma_{\text{stat}}$ error bars; solid line: exact analytical result $I_{2}(\beta)/I_{1}(\beta)$. Biases are $|\Delta|\leq 0.001$ near the training coupling and remain below $0.06$ across $\beta\in[1,4]$.
@@ V.3 Lattice Size Generalisation
Figures 78 illustrate this progression. For the $12\times 8$ lattice (Fig. 7), which shares the training extent $L=8$ in one direction, the model tracks the exact curve well up to $\beta\approx 3$. For the $16\times 16$ lattice (Fig. 8), near-exact agreement is achieved for $\beta\in[1.5,2.5]$, with deviations appearing at higher couplings and at $\beta=1.0$.
Figure 7: Average plaquette $\langle P\rangle$ as a function of coupling $\beta$ for the $12\times 8$ lattice. Blue circles: diffusion model results with $\sigma_{\text{stat}}$ error bars; solid line: exact analytical result $I_{2}(\beta)/I_{1}(\beta)$. The model achieves good agreement across the range $\beta\in[1,4]$ for this geometry, which shares the training extent $L=8$.
Figure 8: Same as Fig. 7 but for the $16\times 16$ lattice. The model achieves near-exact agreement for $\beta\in[1.5,2.5]$.
The plaquette distributions further illustrate the volume dependence. Figure 9 shows the $8\times 12$ distribution at $\beta=2.0$, which remains well-centered on the exact value. In contrast, Fig. 10 for the $32\times 32$ lattice shows a visible shift, with the distribution peaked well below the exact result.
Figure 9: Same as Fig. 5 but for the $8\times 12$ lattice. The distribution remains centered near the exact result, confirming successful generalisation to geometries slightly larger than the training lattice.
Figure 10: Same as Fig. 9 but for the $32\times 32$ lattice at $\beta=2.0$. The distribution is visibly shifted from the exact value (red dashed line), reflecting the difficulty of generalising to lattice sizes much larger than the $8\times 8$ training geometry.
@@ Acknowledgements.
We thank J. Bulava, P. Klein, J. Gegelia, F. He, L. Meng, and W. Sun for valuable discussions and helpful comments on the manuscript.

### v2

@@ I Introduction
The emergence of machine learning approaches offers promising avenues to address these challenges. Normalizing flows have been successfully applied to scalar field theories Albergo2019 and gauge theories Kanwar2020 ; Boyda2021 ; Abbott2023 , demonstrating the potential of generative models in lattice field theory. More recently, diffusion models—a class of generative models that learn to reverse a gradual noising process—have achieved remarkable success in image generation Ho2020 ; Song2021 and have been proposed for lattice gauge theory applications Wang:2023exq ; Wang:2023sry ; Zhu2024a ; Zhu2024b ; Vega:2025hgz ; Kanwar:2025wuc ; Aarts:2026zzr .
@@ IV.4 Training Procedure
The training procedure follows the standard DDPM approach, as illustrated in Fig. 3(a). For each batch of clean configurations $\varphi_{0}$:
@@ IV.5 Sampling Algorithm
To generate new configurations, we reverse the diffusion process starting from pure noise, as illustrated in Fig. 3(b).
Figure 3: Overview of the training and sampling procedures. (a) Flowchart of a single training step. The model learns to predict the noise $\epsilon_{\text{true}}$ that was added to clean configurations, enabling the reverse denoising process during sampling. (b) Sampling algorithm flowchart. Starting from Gaussian noise, the iterative denoising loop applies physics-conditioned noise prediction and gradually recovers a gauge field configuration. The final projection step ensures the output satisfies the SU(2) constraint.
@@ V.1 Validation at Training Coupling
Figure 4(a) shows the plaquette distribution for the training point. The distribution is well-centered on the exact value, with the sample mean nearly coinciding with it, confirming that the model has learned the correct equilibrium distribution at $(\beta_{0},8\times 8)$.
Figure 4: Plaquette distribution from 500 diffusion-generated configurations at $\beta=2.0$. Red dashed line: exact value $I_{2}(\beta)/I_{1}(\beta)$; blue line: sample mean. (a) $8\times 8$ training lattice. The distribution is well-centered on the exact result. (b) $8\times 12$ lattice. The distribution remains centered near the exact result, confirming successful generalisation to geometries slightly larger than the training lattice. (c) $32\times 32$ lattice. The distribution is visibly shifted from the exact value, reflecting the difficulty of generalising to lattice sizes much larger than the $8\times 8$ training geometry.
@@ V.2 Physics-Conditioned Generation at Different Couplings
Figure 5(a) shows the average plaquette as a function of $\beta$. Near the training coupling, the model achieves excellent agreement: at $\beta=1.5$, $2.0$, and $2.5$ the biases are $|\Delta|\leq 0.0004$. At the boundaries of the explored coupling range, larger biases appear: $|\Delta|=0.017$ at $\beta=1.0$ and $|\Delta|=0.051$ at $\beta=3.0$, with intermediate values $|\Delta|\approx 0.01$$0.03$ at $\beta=3.5$ and $4.0$. This demonstrates that physics-conditioned rescaling captures the coupling dependence of the score function well near the training point, with accuracy degrading gradually at larger $|\beta-\beta_{0}|$.
Figure 5: Average plaquette $\langle P\rangle$ as a function of coupling $\beta$. Blue circles: diffusion model results with $\sigma_{\text{stat}}$ error bars; solid line: exact analytical result $I_{2}(\beta)/I_{1}(\beta)$. (a) $8\times 8$ training lattice. Biases are $|\Delta|\leq 0.001$ near the training coupling and remain below $0.06$ across $\beta\in[1,4]$. (b) $12\times 8$ lattice, which shares the training extent $L=8$. The model achieves good agreement across the range $\beta\in[1,4]$. (c) $16\times 16$ lattice. The model achieves near-exact agreement for $\beta\in[1.5,2.5]$.
@@ V.3 Lattice Size Generalisation
Figures 5(b)5(c) illustrate this progression. For the $12\times 8$ lattice (Fig. 5(b)), which shares the training extent $L=8$ in one direction, the model tracks the exact curve well up to $\beta\approx 3$. For the $16\times 16$ lattice (Fig. 5(c)), near-exact agreement is achieved for $\beta\in[1.5,2.5]$, with deviations appearing at higher couplings and at $\beta=1.0$.
The plaquette distributions further illustrate the volume dependence. Figure 4(b) shows the $8\times 12$ distribution at $\beta=2.0$, which remains well-centered on the exact value. In contrast, Fig. 4(c) for the $32\times 32$ lattice shows a visible shift, with the distribution peaked well below the exact result.
@@ Acknowledgements.
We thank J. Bulava, P. Klein, J. Gegelia, F. He, L. Meng, and W. Sun for valuable discussions and helpful comments on the manuscript. The calculations of this work were performed on the HPC cluster Elysium of the Ruhr University Bochum, subsidised by the DFG (INST 213/1055-1).